Q P MC 0 10 ... 1 9 3 2 8 4 3 7 5 4...

Q P MC
0 10 ...
1 9

3

2 8 4
3 7 5
4 6 6
5 5 7

6 . 4 8

The 6, 4, 8 are part of the chart.

The following table shows quantity, price, and marginal cost information for a monopoly. What price should the firm charge to maximize its profit?
Answer #1

Answer

According to profit maximizing condition of a monopoly a firm produces that quantity at which MR = MC

MR = Marginal Revenue is the revenue firm receives by producing 1 addition unit of output

MR for Q = 3 is Total Revenue form 3 units - Total Revenue for 2 units

Total Revenue form 3 units = 7*3 = 21

Total Revenue form 2 units = 2*8 = 16

Hence MR for Q = 3 = 21 - 16 = 5

and MC for Q = 3 is 5

Hence MR = MC for Q = 3

Hence Firm should produce 3 units and Charge Price = 7